# An active learning enabled 4th-generation high dimensional committee neural network potential for the simulation of quantum vibrational spectra: Applications to the bulk and the air-water interface with charge transfer and nuclear quantum effects included

### v1

@@ 4.2 Dipole Moment Validation
For accurate vibrational spectral simulations, one must also assess electronic observables such as dipole-moment fluctuations in addition to energies and forces. We have therefore adopted dipole moments as part of our evaluation criteria. Fig. 3 illustrates dipole moment correlations for the same classical bulk subset of the 2000 configuration validation set described above.
@@
Dipole moment validations for interface and PIMD sets are reported in SI Figs. S3-S5. We obtained dipole moment MAEs of 0.03 D for the classical configurations and 0.04 D for the PIMD configurations across both bulk and interfacial validation sets. These validations demonstrate that the 4G-HDCNNP predicted charges can recover the dipole moment to within a few hundredths of a Debye, proving the accuracy of our employed approach in calculating dipoles. We would like to emphasize that using 4G-HDCNNPs provides a robust and practical paradigm for predictive spectral simulations, eliminating the need for ad-hoc fittings, empirical parameterizations, or charge rescaling to match reference spectra.
Fig. 4 compares the calculated RDFs for bulk water using the final Generation 3 models compared to reference revPBE0-D3.
@@
4G-HDCNNP-generated RDFs from classical MD simulations are in excellent agreement with the reference classical AIMD calculated RDFs for $O_{W}$$O_{W}$, $O_{W}$$H_{W}$, and $H_{W}$$H_{W}$ pairs. This demonstrates that our trained 4G-HDCNNPs can accurately capture the dynamic H-bond network of water at finite temperatures. To evaluate finite-size effects, classical MD simulations were also performed for a larger box of 17.53 Å  containing 180 water molecules, corresponding to the same density of 0.997 g/cm3 at room temperature, where negligible differences were found (see SI Fig. S7).
@@ 4.3 IR Spectra of Bulk Water
The simulated IR spectra of liquid water at 298 K obtained from classical, PA-CMD, and TRPMD simulations using the final trained 4G-HDCNNPs are presented in Fig. 5.
Figure 5: 4G-HDCNNP calculated MD, TRPMD, and PA-CMD simulated IR spectra of bulk water compared to the experiment.52 The dashed line represents the OH-stretch maximum peak position from the experiment.
The experimental IR peak for the O-H stretch region appears at 3420 cm-1. In classical simulations, the O–H stretch exhibits a pronounced blue shift of 274 cm-1 relative to the experiment due to the neglect of NQEs. The blue shift for the classical spectra also appears in the bending region, but to a much lesser extent. The inclusion of NQEs through real-time path integral methods, TRPMD and PA-CMD, largely corrects this deviation. TRPMD generates a stretch peak with a modest blue shift of 24 cm-1 and small broadening effects. The PA-CMD produces a stretch maximum with a slight red shift of 26 cm-1, which can be attributed to the curvature problem commonly associated with PA-CMD. The artificial broadening and red shift of the O-H stretch peak of bulk water at room temperature, as obtained from TRPMD and PA-CMD, respectively, are in agreement with the known problems associated with these methods, as summarized in our previous works.15, 31, 48 All three methods more or less agree on the positions and relative intensities of the peaks for the librational mode, indicating that the 4G models accurately capture the dipole derivatives responsible for low-frequency absorption and can reliably predict underlying vibrational physics.
@@ 4.4 IR Spectra of the Air-Water Interface
Figure 6: Calculated density profile of water in the $z$ direction for the air-water interface system, along with the two considered water layers considered in this work.
Figure 7: 4G-HDCNNP simulated IR spectra of the air-water interface system. Layer 1 corresponds to the bulk-like water at the center of the box, while Layer 2 denotes the surface layer. The deconvoluted spectra are given in the SI Fig. S9.
@@
For the air-water interface, using the density profile of water along the $z$-direction, the simulation box was partitioned into two different layers, namely Layers 1 and 2 (Fig. 6). Layer 1 represents the bulk-like water, which sits close to the center of the box ($z=0$). The thickness of the bulk-like water layer was kept as $\approx$8.00 Å. Layer 2 on each side represents the interface layers with a thickness of $\approx$4.00 Å each. The water density is highest and bulk-like near the center, and shows a gradual depletion toward the vapor phase. Our Layer-resolved IR spectra calculated using classical, TRPMD, and PA-CMD methods for the air-water interface at 298 K are presented in Fig. 7 (the deconvoluted spectra are provided in SI Fig. S9). For each method, the total spectra are accompanied by the spectra from Layers 1 and 2 corresponding to bulk-like and interface-like water molecules, respectively. In the classical simulations, the total spectra exhibit two well-separated peaks. The bulk-like band centered around 3700 cm-1 is dominated by the Layer-1 water molecules, which reside at the center of the box (Fig. 6). The high frequency peak around 3944 cm-1 arises from Layer-2, which is mostly comprised of interfacial water molecules with abundant single H-bond donors and dangling OH bonds. Similar to the bulk spectra, all classical peaks are blue-shifted due to the neglect of the NQEs. The O-H stretch peaks are red-shifted once the NQEs are added through TRPMD and PA-CMD. The TRPMD peak is artificially broadened, which hides the distinct doublet feature characteristic of interfacial systems. The PA-CMD spectrum is expectedly red-shifted and slightly broadened, though to a lesser extent, compared to TRPMD due to its well-known curvature problem. However, the total PA-CMD peak still exhibits a doublet feature, although not as pronounced as the classical spectrum.
4.5 vSFG Spectra of the Air-Water Interface
Figure 8: 4G-HDCNNP simulated vSFG spectra of the air-water interface system. The experimental spectrum is taken from Ref. 53.
vSFG spectra computed using ssVVCF provide a more detailed microscopic picture of the role of NQEs for the considered air-water interface system. Our 4G-HDCNNP-calculated classical, TRPMD, and PA-CMD vSFG spectra are presented in Fig. 8. vSFG exhibits two distinct peaks pointing in opposite directions. The lower-frequency negative lobe corresponds to the H-bonded OH groups, whereas the higher-frequency positive lobe represents the free dangling OH bonds. The experimental spectrum is taken from Ref. 53. We note that our calculated vSFG spectra obtained from all methods do not exhibit the positive features from 3000 cm-1 to 3200 cm-1 in the experimental spectra, in agreement with previous studies.53, 54 The experimental peak for the free OH bonds appears at $\sim$ 3700 cm-1. In our classical spectra, this peak is observed at $\sim$ 3950 cm-1 exhibiting a pronounced $\sim$ 250 cm-1 blue shift. This shift is slightly larger than previously calculated vSFG spectra using dipole–polarizability correlation function and polarizable force fields,55, 56 where the classical free OH peaks show $\sim$ 175 cm-1 blue shift. Similar to our IR spectra, the classical peaks are blue-shifted in the H-bonded region as well. When NQEs are included through TRPMD, both peaks are found to be largely broadened. It should be noted that the ssVVCF method employed here does not account for non-Condon effects, which leads to a shoulder observed at approximately 3650 cm-1 in the experimental free-OH region.54 We anticipate that incorporating non-Condon effects, together with more rigorous vSFG calculations based on the dipole–polarizability correlation function,57 would yield TRPMD spectra in better, more quantitative agreement with the experiment. On the other hand, both the PA-CMD peaks are red-shifted due to the curvature problem, similar to the IR peaks. The PA-CMD-calculated free OH peak is in close agreement with previously reported PA-CMD simulations. 55, 56 These results highlight that 4G-HDCNNPs provide a powerful framework for accurately simulating both IR and vSFG spectra, provided that the underlying quantum dynamic methods and time correlation functions have sufficient accuracy and are free from approximations.
@@ 5 Conclusions
In this work, we have developed, for the first time, 4G-HDCNNPs using active learning for predictive vibrational spectral simulations without the explicit training of dipole moments. The 4G-HDCNNP architecture is combined with path integral molecular dynamics to efficiently incorporate NQEs. Using active learning via QbC, we have shown that 4G-HDCNNPs can be efficiently and accurately trained with only a few hundred configurations. Analyses of IR and vSFG spectra for bulk and air-water interface demonstrate that the 4G-HDCNNPs are reliable. Although the results presented in this work focus on liquid bulk water and its interface with air, we anticipate that the methodology will be broadly applicable to other complex systems and environments with high accuracy. Future work will focus on calculating spectra with higher spectroscopic accuracy through combining 4G-HDCNNPs with recently introduced curvature-free path integral methods branched from CMD, including Te-PIGS,58 f-QCMD,59, 60 and h-CMD15, (ii) extending and applying our 4G-HDCNNP training workflow to complex reactive electrochemical environments at different redox interfaces dominated by non-local charge transfer interactions, and (iii) adding constant potentials to the architecture of 4G-HDCNNPs for the accurate simulations of different interfacial electrochemcial reactions. These developments will enable accurate and predictive vibrational spectral simulations to address important and long-standing questions in condensed-phase and interfacial chemistry.
Acknowledgments
Simulations presented in this work used resources from Bridges-2 at Pittsburgh Supercomputing Center through allocations PHY230030P and CHE250024P from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program,61 which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. The authors also acknowledge the HPC center at UMKC for providing computing resources and support.
@@ Acknowledgments
Details of active learning benchmarks for charge, energy, and force; temperature, energy, and charge convergences from 4G-HDcNNP MD simulations; correlation plots for charge, energy, and force; and calculated RDFs from 4G-HDCNNPs compared to reference rev-PBE0-D3.

### v2

@@ 4.2 Dipole Moment Validation
For accurate vibrational spectral simulations, one must also assess electronic observables such as dipole-moment fluctuations in addition to energies and forces. We have therefore adopted dipole moments as part of our evaluation criteria. Fig. 3 illustrates dipole moment correlations for the same classical bulk subset of the 2000 configuration validation set described above. Dipole moment validations for interface and PIMD sets are reported in the Supporting Information Figs. S3-S5. We obtained dipole moment MAEs of 0.03 D for the classical configurations and 0.04 D for the PIMD configurations across both bulk and interfacial validation sets. These validations demonstrate that the 4G-HDCNNP predicted charges can recover the dipole moment to within a few hundredths of a Debye, proving the accuracy of our employed approach in calculating dipoles. Using 4G-HDCNNPs, therefore, provides a robust and practical paradigm for predictive spectral simulations, eliminating the need for explicit training of dipoles, ad-hoc fittings, empirical parameterizations, or charge rescaling to match reference spectra.
@@
Fig. 4 compares the calculated RDFs for bulk water using the final Generation 3 models compared to reference revPBE0-D3. 4G-HDCNNP-generated RDFs from classical MD simulations are in excellent agreement with the reference classical AIMD calculated RDFs for $O_{W}$$O_{W}$, $O_{W}$$H_{W}$, and $H_{W}$$H_{W}$ pairs. This demonstrates that our trained 4G-HDCNNPs can accurately capture the dynamic H-bond network of water at finite temperatures. To evaluate finite-size effects, classical MD simulations were also performed for a larger box of 17.53 Å  containing 180 water molecules, corresponding to the same density of 0.997 g/cm3 at room temperature, where negligible differences were found (see the Supporting Information Fig. S7).
@@
Figure 5: 4G-HDCNNP calculated MD, TRPMD, and PA-CMD simulated IR spectra of bulk water compared to the experiment.48 The dashed line represents the OH-stretch maximum peak position from the experiment.
Figure 6: Calculated density profile of water in the $z$ direction for the air-water interface system, along with the two water layers considered in this work.
Figure 7: 4G-HDCNNP simulated IR spectra of the air-water interface system. Layer 1 corresponds to the bulk-like water at the center of the box, while Layer 2 denotes the surface layer. The deconvoluted spectra are given in the Supporting Information Fig. S9.
@@ 4.3 IR Spectra of Bulk Water
The simulated IR spectra of liquid bulk water at 298 K obtained from classical, PA-CMD, and TRPMD simulations using the final trained 4G-HDCNNPs are presented in Fig. 5. The experimental IR peak for the O-H stretch region appears at 3420 cm-1. In classical simulations, the O–H stretch exhibits a pronounced blue shift of 274 cm-1 relative to the experiment due to the neglect of NQEs. The blue shift for the classical spectra also appears in the bending region, but to a much lesser extent. The inclusion of NQEs through real-time path integral methods, TRPMD and PA-CMD, largely corrects this deviation. TRPMD generates a stretch peak with a modest blue shift of 24 cm-1 and small broadening effects. The PA-CMD produces a stretch maximum with a slight red shift of 26 cm-1. The artificial broadening and red shift of the O-H stretch peak of bulk water at room temperature, as obtained from TRPMD and PA-CMD, respectively, are in agreement with the known problems associated with these methods, as summarized in our previous works.15, 31, 45 All three methods more or less agree on the positions and relative intensities of the peaks for the librational mode, indicating that the 4G models accurately capture the dipole derivatives responsible for low-frequency absorption and can reliably predict underlying vibrational physics.
@@ 4.4 IR Spectra of the Air-Water Interface
For the IR spectra of the air-water interface, using the density profile of water along the $z$-direction, the simulation box was partitioned into two different layers, namely Layers 1 and 2 (Fig. 6). Layer 1 represents the bulk-like water, which sits close to the center of the box ($z=0$). The thickness of the bulk-like water layer was kept as $\approx$8.00 Å. Layer 2 on each side represents the interface layers with a thickness of $\approx$4.00 Å each. The water density is highest and bulk-like near the center, and shows a gradual depletion toward the vapor phase. Our Layer-resolved IR spectra calculated using classical, TRPMD, and PA-CMD methods for the air-water interface at 298 K are presented in Fig. 7 (the deconvoluted spectra are provided in the Supporting Information Fig. S9). For each method, the total spectra are accompanied by the spectra from Layers 1 and 2 corresponding to bulk-like and interface-like water molecules, respectively. In the classical simulations, the total spectra exhibit two well-separated peaks. The bulk-like band centered around 3700 cm-1 is dominated by the Layer-1 water molecules, which reside at the center of the box (Fig. 6). The high frequency peak around 3944 cm-1 arises from Layer-2, which is mostly comprised of interfacial water molecules with abundant single H-bond donors and dangling free OH bonds. Similar to the bulk spectra, all classical peaks are blue-shifted due to the neglect of the NQEs. The O-H stretch peaks are red-shifted once the NQEs are added through TRPMD and PA-CMD. The TRPMD peak is artificially broadened, which hides the distinct doublet feature characteristic of interfacial systems. The PA-CMD spectrum is expectedly red-shifted and slightly broadened, though to a lesser extent, compared to TRPMD due to its well-known curvature problem. However, the total PA-CMD peak still exhibits a doublet feature, although not as pronounced as the classical spectrum.
@@ 5 Conclusions
In summary, this work demonstrates that 4G-HDCNNPs efficiently trained using active learning and query-by-committee can be utilized for predictive quantum vibrational spectral simulations without the need for explicit training of dipole moments, ad hoc fitting, or parameterizations. This feat is achieved by combining non-local effects from 4G-HDCNNPs with nuclear quantum effects from Feynman’s imaginary-time path integral formalism. Analyses of IR spectra for bulk and air-water interface test cases demonstrate that the 4G-HDCNNP predicted charges are reliable for dipole and IR spectral calculations. The adapted methodology in this work is general and applicable to other complex systems and environments with significant long-range charge transfer effects. Future works will focus on (i) calculating spectra through combining 4G-HDCNNPs with recently introduced curvature-free CMD-derived real-time methods, including Te-PIGS,49 f-QCMD,50, 51 and our very own h-CMD15 methods, (ii) applying 4G-HDCNNPs to complex reactive electrochemical environments at different condensed phases and interfaces dominated by non-local charge transfer interactions, and (iii) including constant potentials for the accurate simulations of different dynamical and spectral properties in the presence of external electric fields. We hope that these developments will facilitate the accurate calculation of quantum dynamical properties and vibrational spectra, addressing important and long-standing questions at the heart of energy conversion and storage processes, among others.
@@
The data that support the findings of this study are provided in the main text and the accompanying Supporting Information. Additional data are available from the corresponding author upon request.
{acknowledgement}
Simulations presented in this work used resources from Bridges-2 at Pittsburgh Supercomputing Center through allocations PHY230030P and CHE250024P from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program,52 which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. The authors also acknowledge the HPC center at UMKC for providing computing resources and support.